OpenStudy (anonymous):

The table shows the total cost of cupcakes (y) for (x) number of cupcakes. Number of Cupcakes Total Cost x y 3 6 4 8 5 10 6 12 Part a. Write a function rule to relate (y) and (x). Part b. Show how you can use the function rule to find the total cost for 11 cupcakes. Show all of your work

OpenStudy (anonymous):

3*2=6, 4*2=8, 5*2=10, 6*2=12, as you can see, y is found by multiplying x by 2. So the function rule to relate y and x is- y=2x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Do you know how to find part B?

OpenStudy (anonymous):

Just plug in 11 for x in y=2x. So it would be y=2(11)

OpenStudy (anonymous):

And simplify by multiplying 2 and 11

OpenStudy (anonymous):

o ok thank you